OpenStudy (anonymous):

Look at the system of equations shown below. (a) 3x - 4y = -1 (b) 6x – 3y = 3 Which could be the first steps to solve this system of equations by elimination? A. Multiply equation (a) by 2, then add the two equations. B. Multiply equation (a) by -2, then add the two equations. C. Multiply equation (b) by 4, then add the two equations. D. Multiply equation (b) by -4, then add the two equations.

OpenStudy (anonymous):

letter B (a) by -2

OpenStudy (anonymous):

How'd you come out with A? @nathan_czh How'd you get B? @bahupepe

OpenStudy (anonymous):

B. Multiply equation (a) by -2, then add the two equations.

OpenStudy (anonymous):



OpenStudy (anonymous):

A. Multiply equation (a) by 2, then add the two equations. **IS NOT RIGHT**

OpenStudy (anonymous):

-2 (3x - 4y = -1) -2 + 6x – 3y = 3 -6x+8y=2 + 6x – 3y = 3 6x will be cancel, and the result is 5y=5 threfore y=1

OpenStudy (anonymous):

you have to multiply (a) by -2 so that when you add them, the -6x and the +6x cancel out.

OpenStudy (anonymous):

Whoops. Didn't see the add